# A Survey of Threats Against Voice Authentication and Anti-Spoofing Systems

### v1

@@ 5 Adversarial Attacks
Adversarial attacks involve subtle, often hard-to-notice changes to speech signals designed to mislead SRS. Unlike spoofing, adversarial attacks directly alter the model’s decision boundaries. As speaker recognition technologies have evolved from traditional models, like GMM-UBM pipelines, to deep learning-based models like x-vector [37] and ECAPA-TDNN [60], the complexity and attack realism strategies have also grown. These strategies have evolved from purely digital changes to real-world, over-the-air implementations. Table. 3 presents a summary of adversarial attacks.
@@ 5.1 Foundational Work
The threat of adversarial audio was first examined by Hautamäki et al. in their study [61]. They employed professional impersonators to target SV systems. Although false acceptance rates rose slightly, from 9.03% to 11.61%, these mimicry attacks showed that human-level imitation is not a serious danger when compared to synthetic methods.
A key change occurred with Kreuk et al. [62], who used the Fast Gradient Sign Method (FGSM) to create targeted adversarial changes for end-to-end neural models. This significantly reduced accuracy from 85.5% to 37.5% on YOHO [63]) and dramatically increased false-positive rates. Additionally, their adversarial examples transferred between datasets and architectures, providing strong evidence of deep vulnerabilities in neural SRS.
@@ 5.2 White-Box Attacks
Li et al. [64, 65] extended these attacks to traditional systems such as GMM-UBM and i-vectors. By applying FGSM to the feature space, they raised false acceptance rates to nearly 100% under imperceptible perturbations. In later work, they employed room impulse response (RIR) modeling to simulate real acoustic environments, achieving up to 50% EER in black-box transfer scenarios, moving adversarial attacks closer to real-world feasibility.
Joshi et al. [66] analyzed modern x-vector-based systems under white-box conditions, showing that attacks like BIM, PGD, and Carlini-Wagner reduced classification accuracy to near-zero even under minimal perturbation ($L_{\infty}=0.001$). These results underscore the inherent fragility of deep speaker representations in the absence of robust defenses.
@@ 5.3 Black-Box Attacks
More recent efforts focus on realistic attack settings with limited model access. Zhang et al. [67] introduced VMask, which uses psychoacoustic masking to guide perturbation generation, significantly boosting false acceptance (e.g., 3.9% to 88.2%) while preserving perceptual quality.
@@
FakeBob [22] demonstrated a black-box attack using Natural Evolution Strategy (NES) and BIM to achieve over 90% targeted success on i-vector systems. However, the method required thousands of queries per sample and prior knowledge of the target model, limiting its practicality for real-time deployment.
Chen et al.  [68] proposed QFA2SR, that pushed further by developing a query-free attack that uses ensemble-based surrogate modeling and knowledge distillation. The attack reached 89.3% success on commercial APIs like Microsoft Azure, though it demanded significant computational resources due to surrogate training and optimization complexity.
5.4 Physical Attacks
@@
To bridge the gap between simulation and real-world application, Chen et al.  [69] proposed a generalizable, transfer-based framework that supports SV, CSI, and OSI. Using multi-task learning and RIR simulation, AS2T achieved success rates up to 98.5% in white-box settings, demonstrating cross-task generalization.
@@
PhyTalker [70] brought adversarial attacks into real-time physical space by injecting phoneme-level perturbations during live playback via speakers. Achieving 85% targeted success in real-world conditions without model access, PhyTalker showed how physical attacks can defeat even robust systems without retraining or dataset access.
Complementarily, Kassis et al. [71] used pitch and duration manipulations to clone prosodic patterns, bypassing both speaker verification and spoofing defenses. Although effective (¿90% ASR), this method required over 100 high-quality utterances per target speaker, limiting its scalability.
Table 3: Comparison of Key Adversarial Attacks on Voice Authentication Systems.
@@
Study Year Attack Target SRS Method Results Limitations / Notes
@@
[61] 2013 Human mimicry GMM-UBM, i-vector Professional impersonation FAR ↑ 9.03% → 11.61% Less effective vs. synthetic attacks
[62] 2018 White-box, transferable End-to-end (YOHO, NTIMIT) FGSM on MFCC Acc ↓ 85.5% → 37.5%, FPR ↑ 4.88% → 94.63% Strong transferability
[64] 2020 White-box, Black-box GMM-UBM, x-vector FGSM + RIR sim. FAR ↑ to 99.99%, EER ↑ to 50% Imperceptible noise; transferable
[66] 2021 White-box x-vector BIM, PGD, CW attack Acc ↓ to  0% at $L_{\infty}$=0.001 Extremely vulnerable even with small $\ell_{\infty}$
[69] 2022 White-box, transferable SV, CSI, OSI Multi-task + RIR sim. Acc ↑ 98.5% (SV), 97.6% (CSI) Strong generalization
[67] 2020 Black-box (stealth) i-vector, x-vector Psychoacoustic masking (VMask) FAR ↑ 3.9% → 88.2% High stealth, indistinguishable audio
[22] 2021 Black-box (targeted) GMM-UBM, i-vector NES + BIM ASR: 90.5% (i-vector), 78.6% (GMM-UBM) High query cost/sample
[68] 2023 Black-box (query-free) Azure API Surrogate ensemble + KD ASR ↑ to 89.3% Costly surrogate training
@@
[70] 2022 Black-box (physical) Real-time SRS Time-domain phoneme injection ASR ↑ 85% OTA Real-time; no model access
[71] 2023 Physical (signal) SRS + Anti-spoof Prosody + spectral cloning ASR ↑ >>90% (ASVspoof, SASV) Needs 100+ target utterances
@@ 6 Deepfake Attacks
Deepfake speech attacks exploit advanced voice synthesis techniques, such as VC and TTS, to impersonate users in SRS. These attacks can bypass both automated authentication systems and human listeners by generating speech that closely mimics a target speaker’s voice, often using minimal training data. Unlike traditional spoofing, which relies on replayed audio, deepfake attacks can fabricate entirely new utterances, expanding the threat surface considerably. As generative models become more accessible and natural-sounding, deepfake attacks pose a growing challenge to voice-based security systems. A summary of these attacks is illustrated in Table. 4.
Scope Note: Although a large body of work exists on high-quality speech synthesis and voice cloning technologies—such as Tacotron 2 [72], VITS [73], and Vall-E [74]—we limit our review to studies that explicitly evaluate these technologies in the context of adversarial attacks against speaker recognition systems. General-purpose TTS or voice conversion models not assessed in a threat setting are excluded to maintain focus on empirical evidence of system vulnerabilities [75, 76, 72, 77, 74, 3, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 57, 96, 97].
6.1 Foundation Work
One of the earliest empirical demonstrations of AI-powered voice spoofing was conducted by Saxena et al. [98]. Using an off-the-shelf voice morphing tool, they reconstructed a target speaker’s voice using only a few minutes of enrollment data. The manipulated voices were evaluated against both automated speaker verification systems and human listeners. Their results were striking: state-of-the-art SRS failed to reject 80–90% of spoofed attempts, and human listeners only correctly identified fakes about 50% of the time, particularly when the attacker mimicked familiar individuals.
This study served as an early warning about the potency of voice conversion attacks and underscored the limitations of both machine-based and human auditory verification. It also emphasized the urgency of developing liveness detection and anti-spoofing strategies before high-fidelity synthesis technologies became more widespread.
@@ 6.2 Partial Fake (PF) Attacks
More recent work has shifted toward PF speech, where only segments of an utterance are synthetically generated. Alali et al. [99] explored PF attacks using zero-shot TTS and VC models, such as VITS and YourTTS. Their findings reveal that both humans and SRS models remain highly susceptible to these nuanced manipulations. Partial fakes—because they preserve some genuine speech—are often more difficult to detect than fully synthetic samples, especially in zero-shot scenarios where no speaker-specific fine-tuning is needed.
Despite demonstrating strong attack success across multiple systems, the study leaves several open questions. It lacks evaluation under noisy or real-world deployment conditions (e.g., background speech, channel variation) and does not assess defense mechanisms or adversarial robustness. These limitations reduce the actionable value of the findings for real-time or production-grade systems.
Table 4: Summary of Deepfake Attacks on Voice Authentication Systems.
Study Year Type Target Method Effect Limitations
[98] 2015 Full Voice Imitation ASV & Human Voice morphing from a few minutes of speech ASV bypass ↑ (80–90%), Human error → 50% Early tools; no env. or cross-lingual tests
[99] 2025 Partial Fake Speech SRS & Human Zero-shot TTS (YourTTS), VC (VITS) for partial synthesis Detection ↓ (systems & humans) Noisy/multilingual settings untested; defense eval. limited
@@ 7 Adversarial Spoofing Attacks
Anti-spoofing countermeasures (CMs), also known as presentation attack detection (PAD), are a critical defense line in SV pipelines, designed to distinguish bona fide speech from spoofed audio. However, recent studies have demonstrated that these systems are highly vulnerable to adversarial attacks, which introduce crafted perturbations to degrade the reliability of CMs and increase the false acceptance of synthetic or replayed speech. Table. 5 presents a summary of these attacks.
@@ 7.1 White-Box Attacks
Early efforts, by Liu et al. [100], utilized white-box adversarial techniques, such as FGSM and PGD, to PAD models. Even minimal perturbations ($\epsilon$ below human perceptibility) caused detector error rates to rise drastically—up to 85–95% in some cases. These results reveal that, with access to model gradients, attackers can nearly eliminate the PAD’s ability to detect spoofed speech, effectively reducing its performance to random guessing.
Building on this, Gomez-Alanis et al. [101] proposed a trainable adversarial generator named ABTN (Adversarial Biometrics Transformation Network), designed to optimize perturbations using PAD and ASV losses. ABTN outperformed traditional FGSM by generating perturbations that generalized better across spoof types and preserved the speaker’s identity, thus fooling PAD without raising flags in the ASV system. It improved attack EER by 27–51% (absolute) across ASVspoof 2019 logical and physical access protocols [102].
@@ 7.2 Black-Box Attacks
To overcome the limitations of white-box assumptions, several works have explored black-box strategies. Zhang et al. [67] developed VMask, which utilizes psychoacoustic masking to craft perturbations imperceptible to humans but effective against commercial and open-source systems. However, its real-world applicability was hindered by playback sensitivity and environmental noise.
GANBA [103] introduced a conditional GAN architecture capable of simultaneously fooling both the speaker verification and PAD modules. Trained on bona fide and spoofed samples, the model generated synthetic utterances with high perceptual quality. Evaluations on ASVspoof 2019 showed that speaker verification accuracy dropped below 20%, and PAD performance significantly degraded, illustrating the danger of jointly optimized, high-quality adversarial audio.
A more physically robust approach was introduced by Kassis et al. [71], who proposed a lightweight signal processing attack that successfully fooled both ASV and CM systems under over-the-air playback. By cloning prosodic and spectral features from high-quality speaker recordings, the attack bypassed multiple detection systems; however, it required at least 15 minutes of speech from the target speaker and was less effective against advanced models, like Wav2Vec-based PADs.
Xuan Hai et al. [104] presented SiFDetectCracker, a black-box query-based attack that perturbs non-speech segments (e.g., silence, background noise) to exploit the reliance of PADs on environmental cues. The attack achieved over 80% success rate without altering speaker-identifying components, thereby remaining effective for ASV and human listeners. This highlights that adversarial success does not require manipulating the entire signal—targeting peripheral acoustic features may suffice.
Zuo et al. [105] presented AdvTTS, a two-stage framework that combines zero-shot TTS with PGD-based perturbation of speaker embeddings. The attack achieved strong results in black-box settings. However, relies on surrogate models and incurs some degradation in naturalness, which may impact stealth in real-world applications.
A different attack type involves creating universal adversarial filters. Panariello et al. [106] introduced Malafide, a fixed finite impulse response (FIR) filter learned via adversarial training. Once optimized, the filter can be convolved with any spoofed utterance to evade detection, without needing input-specific perturbation. The method was effective even against top-performing deep PADs and is practical for real-time application via playback devices or frequency response shaping.
This approach was extended in Malacopula [107], a nonlinear filter targeting the ASV component itself by enhancing spoofed audio to better match the victim’s voice. These methods demonstrate a trend toward physical-world-capable adversarial strategies that survive digital-to-analog transformations, re-recording, and transmission through real-world channels.
@@
Study Year Access Method Target Effect Notes
[100] 2019 White-box Gradient-based perturbations PAD $\uparrow$ Error rate >>85% (large $\epsilon$) Needs gradients; small-norm bound
[101] 2022 White-box ABTN (adversarial generator, PAD+ASV loss) PAD, ASV $\uparrow$ Attack EER +27–51% Generalizes better than FGSM; identity preserved
[103] 2022 White-box GAN-generated spoof speech PAD, ASV $\downarrow$ ASV Acc <<20%; PAD degraded High compute; training overhead
[67] 2020 Black-box Psychoacoustic, mimicry PAD $\uparrow$ ASR; realistic audio Playback constraints; noise-sensitive
[71] 2023 Black-box Playback perturbation ASV, CM $\downarrow$ Acc even under replay Needs 15+ min data; weak on strong CMs
[104] 2023 Black-box Perturbing background/silence via score queries PAD $\uparrow$ >>80% ASR; identity preserved Query-heavy; PAD-specific
[106] 2023 Black-box (univ.) FIR adversarial filter PAD $\rightarrow$ Universal spoofing; real-time feasible Attack-tuned; generalization limited
[107] 2024 Black-box (univ.) Nonlinear adversarial filter (Malacopula) ASV $\uparrow$ Voice match; bypasses ASV Targets ASV instead of PAD; physical-world viable
[105] 2024 Black-box Zero-shot TTS, PGD on speaker emb. SI, PAD $\uparrow$ Transferable; bypasses ADD Uses surrogate models; quality drop

### v2

@@ 5 Adversarial Attacks
Adversarial attacks introduce subtle, often imperceptible, perturbations into speech signals to mislead Speaker Recognition Systems (SRS). Unlike spoofing attacks, which typically involve synthesized or replayed speech, adversarial attacks directly manipulate the input data to exploit and alter the model’s decision boundaries. As speaker recognition technology has transitioned from traditional methods, such as Gaussian Mixture Model-Universal Background Model (GMM-UBM) pipelines, to deep learning-based architectures like x-vectors [37] and ECAPA-TDNN [60], the sophistication of attack strategies has advanced in parallel, evolving from purely digital manipulations to practical over-the-air implementations. Table 3 summarizes key adversarial attacks on speaker recognition systems.
@@ 5.1 Foundational Work
The concept of adversarial audio in speaker verification was first explored by Hautamäki et al. [61], who employed professional impersonators to target Speaker Verification (SV) systems. Although the false acceptance rate (FAR) increased only slightly from 9.03% to 11.61%, these mimicry attacks demonstrated that human impersonation poses a less significant threat than computational, synthetic methods.
A pivotal shift occurred with the work of Kreuk et al. [62], who applied the Fast Gradient Sign Method (FGSM) to generate targeted adversarial perturbations for end-to-end neural models. This attack reduced accuracy from 85.5% to 37.5% on the YOHO dataset [63] and increased the false positive rate from 4.88% to 94.63%. Furthermore, the adversarial examples demonstrated transferability across different datasets and model architectures, highlighting a systemic vulnerability in neural SRS.
@@ 5.2 White-Box Attacks
Li et al. [64] first extended these gradient-based attacks to traditional systems, including GMM-UBM and i-vector models. By applying FGSM to the feature space, they increased the FAR to nearly 100% using imperceptible perturbations. Building on this, Li et al. [65] designed practical attacks against the state-of-the-art x-vector system, with both untargeted and targeted objectives. A key innovation was the integration of an estimated Room Impulse Response (RIR) to craft perturbations robust enough to withstand real-world playback distortions. The approach achieved a 98% attack success rate (ASR) in digital simulations and 50% in over-the-air physical playback tests, marking a significant advance in physically realizable adversarial attacks.
Further exploring these vulnerabilities, Joshi et al. [66] analyzed modern x-vector-based systems under white-box conditions. They demonstrated that iterative attacks such as the Basic Iterative Method (BIM), Projected Gradient Descent (PGD), and Carlini-Wagner (CW) could reduce classification accuracy to nearly zero, even with a minimal perturbation budget ($\ell_{\infty}\leq 0.001$). In parallel, Zhang et al. [67] developed VMask, which utilizes psychoacoustic masking to craft perturbations that are imperceptible to humans but effective against both commercial and open-source systems. However, its real-world applicability was hindered by sensitivity to playback hardware and environmental noise. These findings collectively underscore the fragility of deep speaker representations against a variety of white-box threats.
@@ 5.3 Black-Box Attacks
The FakeBob attack [22] demonstrated a practical black-box approach by combining the Natural Evolution Strategy (NES) and BIM to achieve a targeted ASR exceeding 90% against i-vector systems. However, its practicality was limited by its high query cost, requiring thousands of model queries per sample.
To address the high query requirement, Chen et al. [68] proposed QFA2SR, a query-free black-box attack. It leverages ensemble-based surrogate models and knowledge distillation to generate adversarial examples without direct access to the target system. This method achieved an 89.3% ASR against commercial APIs, including Microsoft Azure, but its reliance on complex surrogate model training requires significant computational resources.
5.4 Physical Over-the-Air Attacks
@@
To bridge the gap between digital simulations and real-world applications, Chen et al. [69] proposed AS2T, a generalizable transfer-based framework targeting SV, CSI, and OSI. By utilizing multi-task learning and RIR simulation, AS2T achieved an ASR of up to 98.5% in white-box settings and demonstrated strong cross-task generalization.
@@
PhyTalker [70] extended adversarial attacks to the real-time physical domain by injecting precisely timed, phoneme-level perturbations during live playback. PhyTalker achieved an 85% targeted ASR in over-the-air tests without requiring access to the target model’s parameters, demonstrating that physical attacks can compromise robust systems in real-time.
Table 3: Comparison of Key Adversarial Attacks on Speaker Recognition Systems.
@@
Reference Year Attack Type Target System Method Key Results Key Findings / Limitations
@@
Hautamäki et al. [61] 2013 Human Mimicry GMM-UBM, i-vector Professional impersonation FAR: 9.03% $\rightarrow$ 11.61% Less effective than computational synthetic methods.
White-Box Attacks
Kreuk et al. [62] 2018 Gradient-based, Transferable End-to-end Neural Nets FGSM on MFCCs Accuracy: 85.5% $\rightarrow$ 37.5%
FPR: 4.88% $\rightarrow$ 94.63%
Demonstrated high transferability across models and datasets.
Li et al. [64] 2020 Gradient-based, Transferable GMM-UBM, i-vector FGSM on feature space FAR $\approx$ 100% (white-box)
EER up to 50% (black-box)
Achieved high success with imperceptible noise; transferable.
Li et al. [65] 2020 Over-the-Air (OTA) x-vector Gradient optimization + RIR estimation ASR: 98% (digital), 50% (OTA) Focused on practical, physically robust perturbations.
Joshi et al. [66] 2021 Iterative Gradient-based x-vector BIM, PGD, CW attacks Accuracy $\rightarrow$ 0% at $\ell_{\infty}\leq 0.001$ Showed extreme vulnerability to minimal perturbations.
Black-Box Attacks
Zhang et al. [67] 2020 Score-based, Stealth i-vector, x-vector Psychoacoustic masking (VMask) FAR: 3.9% $\rightarrow$ 88.2% Perturbations are psychoacoustically masked to be imperceptible.
Chen et al. [22] 2021 Score-based, Targeted (FakeBob) GMM-UBM, i-vector Natural Evolution Strategy + BIM ASR: 90.5% (i-vector) Effective but requires thousands of queries per sample.
Chen et al. [68] 2023 Query-free, Transfer-based (QFA2SR) Commercial APIs (Azure) Surrogate ensemble + Knowledge Distillation ASR up to 89.3% Eliminates query cost but requires costly surrogate model training.
Physical Attacks
Chen et al. [69] 2022 Transfer-based (AS2T) SV, CSI, OSI systems Multi-task learning + RIR simulation ASR up to 98.5% Demonstrates strong generalization across different speaker recognition tasks.
@@
Chen et al. [70] 2022 Real-time (PhyTalker) Real-time SRS Time-domain phoneme injection Targeted ASR: 85% (OTA) Real-time physical attack without requiring model access.
@@ 6 Deepfake Attacks
Deepfake speech attacks exploit advanced voice synthesis techniques, such as Voice Conversion (VC) and Text-to-Speech (TTS), to impersonate users in Speaker Recognition Systems (SRS). These attacks can bypass both automated authentication systems and human listeners by generating speech that closely mimics a target speaker’s voice, often using minimal training data. Unlike traditional replay spoofing, deepfake attacks can fabricate entirely new utterances, significantly expanding the threat landscape. As generative models become more powerful and accessible, these attacks pose a growing challenge to voice-based security. A summary of key deepfake attacks is presented in Table 3.
Scope Note: This review focuses specifically on studies that evaluate speech synthesis technologies as attack vectors against speaker recognition systems. While the field of generative audio synthesis is vast, encompassing foundational models like Tacotron 2 [71] and advanced architectures like VALL-E [72], we exclude works that do not provide an empirical assessment of their models in an adversarial context. This scope ensures our analysis remains centered on demonstrated vulnerabilities in voice biometrics. For a broader overview of modern speech synthesis technologies, see [73, 74, 71, 75, 72, 3, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 57, 94, 95].
6.1 Foundational Work
One of the earliest empirical demonstrations of AI-driven voice spoofing was presented by Saxena et al. [96]. Using a commercially available voice morphing tool, they synthesized a target speaker’s voice from just a few minutes of enrollment data. The resulting audio was evaluated against both automated speaker verification systems and human listeners. The results were striking: state-of-the-art SRS accepted 80–90% of the spoofed attempts, while human listeners correctly identified the fakes only about half the time.
This study served as an early warning of the potency of voice conversion attacks and underscored the limitations of both machine-based and human auditory verification. It also highlighted the urgent need for robust anti-spoofing and liveness detection mechanisms as high-fidelity synthesis technologies became more widespread.
@@ 6.2 Partial Fake (PF) Attacks
Recent research has explored the threat of partial fake (PF) speech, where only specific segments of an utterance are synthetically generated while others remain authentic. Alali et al. [97] investigated PF attacks using modern zero-shot TTS and VC models, including VITS and YourTTS. Their findings revealed that both humans and SRS models are highly susceptible to these nuanced manipulations. Because they preserve characteristics of genuine speech, partial fakes are often more challenging to detect than fully synthetic samples, particularly in zero-shot scenarios that require no speaker-specific fine-tuning.
Although the study demonstrates high attack success rates, it also highlights several research gaps. The attacks were not evaluated under noisy or other challenging real-world conditions, such as those involving channel variability or background speech. Consequently, the robustness of these attacks and potential countermeasures in practical deployment scenarios remains an open area for investigation.
Table 4: Summary of Key Deepfake Attacks on Voice Authentication Systems.
Reference Year Attack Type Target Method Key Results Limitations / Notes
Saxena et al. [96] 2015 Full Voice Conversion ASV & Human Listeners Voice morphing using a few minutes of source audio ASV bypass rate: 80–90%
Human accuracy $\approx$ 50%
Used early-generation tools; lacked environmental or cross-lingual testing.
Alali et al. [97] 2025 Partial Fake Speech SRS & Human Listeners Zero-shot TTS (YourTTS) and VC (VITS) for partial synthesis High success rates; reduced detection accuracy for both systems and humans. Not evaluated in noisy or multilingual settings; defenses not thoroughly assessed.
@@ 7 Adversarial Spoofing Attacks
Anti-spoofing countermeasures (CMs), also known as Presentation Attack Detection (PAD), are a critical defense layer in Speaker Verification (SV) pipelines, designed to distinguish bona fide speech from spoofed audio. However, recent studies have demonstrated that these systems are themselves highly vulnerable to adversarial attacks. These attacks introduce crafted, often imperceptible, perturbations to degrade the reliability of CMs, thereby increasing the false acceptance of synthetic or replayed speech. Table 5 presents a summary of these attacks.
@@ 7.1 White-Box Attacks
Early work by Liu et al. [98] applied white-box adversarial techniques, such as the Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD), against PAD models. Even minimal, imperceptible perturbations (i.e., small $\epsilon$) increased detector error rates to as high as 85–95%. These results revealed that, with access to model gradients, an attacker can nearly eliminate a PAD system’s ability to detect spoofed speech, effectively reducing its performance to that of random guessing.
Building on this, Gomez-Alanis et al. [99] proposed a trainable adversarial generator named ABTN (Adversarial Biometrics Transformation Network), designed to optimize perturbations using both PAD and Automated Speaker Verification (ASV) losses. ABTN outperformed traditional FGSM by generating perturbations that generalized better across spoof types while preserving the speaker’s identity. This allowed the attack to fool the PAD system without compromising the ASV system’s ability to verify the speaker. The method improved the attack Equal Error Rate (EER) by an absolute 27–51% on the ASVspoof 2019 logical and physical access protocols [100].
A different approach, GANBA [101], introduced a conditional Generative Adversarial Network (GAN) architecture capable of simultaneously fooling both the SV and PAD modules. Trained on bona fide and spoofed samples, the model generated high-quality synthetic utterances that significantly degraded system performance. Evaluations on ASVspoof 2019 showed that SV accuracy dropped below 20% while PAD effectiveness was severely compromised, illustrating the threat posed by jointly optimized adversarial audio.
@@ 7.2 Black-Box Attacks
To overcome the limitations of white-box assumptions, several studies have explored black-box strategies that do not require gradient access.
Xuan Hai et al. [102] presented SiFDetectCracker, a query-based black-box attack that perturbs non-speech segments (e.g., silence, background noise) to exploit the reliance of some PAD systems on environmental cues. The attack achieved a success rate of over 80% without altering speaker-identifying vocal components, ensuring the audio was still accepted by the ASV system. This highlights that adversarial success does not require manipulating the entire signal—targeting peripheral acoustic features can be sufficient.
A more physically robust approach was introduced by Kassis et al. [103], who proposed a lightweight signal processing attack that successfully deceived both ASV and CM systems during over-the-air playback. By cloning prosodic and spectral features from high-quality speaker recordings, the attack bypassed multiple detection systems. However, it was less effective against advanced models, such as Wav2Vec2-based PADs. The attack also assumes the attacker can inject audio via a rooted phone, a prerequisite often blocked by the security measures of high-stakes applications.
Another category of black-box attacks involves creating universal adversarial filters. Panariello et al. [104] introduced Malafide, a fixed Finite Impulse Response (FIR) filter learned via adversarial training. This universal filter can be convolved with any spoofed utterance to evade detection, removing the need to craft input-specific perturbations. This method proved effective against high-performing deep PADs and is practical for real-time applications.
This approach was extended in Malacopula [105], which created a nonlinear filter targeting the ASV component itself by enhancing spoofed audio to better match the victim’s voiceprint. These methods demonstrate a clear trend toward physically realizable adversarial strategies that can survive digital-to-analog conversion and transmission through real-world channels.
@@
Reference Year Access Method Target Effect on Target Notes / Limitations
White-Box Attacks
Liu et al. [98] 2019 White-box Gradient-based perturbations (FGSM, PGD) PAD Increases error rate to >>85% Requires model gradient access; bounded by small perturbation norm.
Gomez-Alanis et al. [99] 2022 White-box ABTN (Adversarial generator with PAD+ASV loss) PAD, ASV Increases attack EER by 27–51% (absolute) Generalizes better than FGSM while preserving speaker identity.
Gomez-Alanis et al. [101] 2022 White-box GAN-based spoofed speech generation (GANBA) PAD, ASV Reduces ASV accuracy to <<20%; degrades PAD Computationally intensive; requires extensive training.
Black-Box Attacks
Hai et al. [102] 2023 Black-box Query-based perturbation of background/silence PAD Achieves >>80% ASR while preserving speaker identity Attack is query-heavy and tailored to specific PAD vulnerabilities.
Kassis et al. [103] 2023 Black-box Over-the-air signal processing attack ASV, CM Bypasses multiple systems via physical playback Requires >>15 min of target data; less effective on robust CMs.
Panariello et al. [104] 2023 Black-box (Universal) Universal FIR adversarial filter (Malafide) PAD Enables real-time, universal bypass of a target PAD Filter is universal for one PAD but may not generalize to others.
Todisco et al. [105] 2024 Black-box (Universal) Universal nonlinear filter (Malacopula) ASV Enhances voice match of spoofed audio to bypass ASV Targets the ASV system directly, not the PAD; physically viable.